OpenStudy (anonymous):

If h(x) = 3x − 1 and j(x) = −2x, find h[j(2)]

OpenStudy (anonymous):

If f(x) = −3x − 3 and g(x) = 2x + 1, find (−2).

OpenStudy (anonymous):

If f(x) = x − 2 and g(x) = −2x − 1, find (f • g)(9)